CaCO3(s) ⇄ CaO(s) + CO2(g)

0.100 mol of CaCO3 and 0.100 mol CaO are placed in an 10.0 L evacuated container and heated to 385 K. When equilibrium is reached the pressure of CO2 is 0.220 atm. 0.300 atm of CO2 is added, while keeping the temperature constant and the system is allowed to reach again equilibrium. What will be the final mass of CaCO3?

7.47 g

12.54 g

18.01 g

2.00 g

10.01 g
Answer #1

Answer:

Step 1: write the balanced chemical equation

CaCO3(s) <-------------> CaO(s) + CO2(g)

Step 2: Calculate the amount before adding any additional CO2

K = Pco2    [ all other elements are in solid condition so the value will be 1 ]
K = 0.220 atm

By Using  ideal gas formula, we have

n = PV/RT = ( 0.220 atm ×10 L ) / (0.08206 L-atm mol-1 K-1 × 385 K) = 0.0696 mol

This is the amount of CaCO3 which has been converted to CaO before pumping-in additional 0.300 atm CO2(g).

Step 3: Calculate the amount after adding the additional CO2

After pumping-in additional 0.300 atm CO2(g), the equilibrium CO2 pressure is still 0.220 atm . Therefore all this additional CO2 would completely convert to CaCO3.

n = PV/RT = ( 0.300 atm ×10 L ) / (0.08206 L-atm mol-1 K-1 × 385 K) = 0.0949 mol

Hence the total CaCO3 after equilibrium is re-established is:

0.100 mol - 0.0696 mol + 0.0949 mol = 0. 1253 mol  

Step 4: calculate the final mass of CaCO3

Now convert mole into mass of CaCO3

We know, mass = moles × molar mass of CaCO3 = 0.1253 mol × 100.0869 g/mol = 12.54 g